# Cardinality-Preserving Structured Sparse Graph Transformers for Molecular Property Prediction

### v2

@@ Abstract
Drug discovery motivates accurate molecular property prediction when labeled data are limited and candidate spaces are vast. This article presents CardinalGraphFormer, a graph transformer that combines structured sparse attention with graph-specific structural biases and an unnormalized aggregation channel that preserves neighborhood cardinality. The model restricts attention to local shortest-path neighborhoods for efficiency, injects degree-based centrality and direct-bond edge information as biases, and augments normalized attention with a gated sum over the same support set to retain dynamic neighborhood-size signals. Pretraining unifies two self-supervised objectives: contrastive alignment of two augmented graph views and masked reconstruction of node and edge attributes. We evaluate with a matched protocol across multiple public molecular property benchmarks and compare against objective-matched transformer baselines as well as masked-modeling and contrastive baselines. Results show consistent improvements attributable to the cardinality-preserving channel and dual pretraining, alongside ablation evidence and efficiency measurements on large molecules. The work emphasizes reproducibility through shared architectures, data filtering, and split protocols, and discusses limitations tied to sparsity benefits on small graphs.
Keywords: graph transformers; molecular property prediction; self-supervised learning; structured sparsity; cardinality-preserving attention.
@@ 1 Introduction
Molecular property prediction serves as a key proxy task in drug discovery, constrained by limited experimental labels and an immense candidate space [6, 1]. Graph transformers with self-supervised pretraining have shown strong performance on standardized benchmarks [9, 10, 3]. We adopt Graphormer-inspired structural biases (shortest-path distance (SPD) + centrality; direct-bond edge bias) but implement structured sparse attention using $K$-hop SPD masks for locality and efficiency. We further augment this with a cardinality-preserving attention (CPA) channel to retain dynamic support-size signals complementary to static centrality embeddings.
@@
  • A structured sparse graph transformer layer with aligned cardinality-preserving attention.
@@
  • Rigorous matched-protocol evaluation showing consistent, often significant gains over reproduced baselines. Code and artifacts are provided in the repository; a stable release commit will be frozen upon acceptance.
@@ 3.2 Structured Sparse Attention with Graphormer-Inspired Biases
The support set is $\mathcal{S}(i)=\{j\mid\mathrm{SPD}(i,j)\leq K\}$ (includes self, SPD(i,i)=0(i,i)=0). Main experiments use $K=3$. Attention is masked to $\mathcal{S}(i)$ (outside: -$\infty$). Logits:
@@
where $\phi(d)=\min(d,K)$ yields $K+1$ bins $\{0,\dots,K\}$, $b_{e_{ij}}$ is a learnable embedding of direct-bond features (bond type, aromaticity, conjugation, ring membership, and stereo configuration; categorical fields are one-hot encoded and embedded separately, then summed; stereo uses an explicit ‘none’ category when unspecified or absent; edge features are symmetric for the undirected graph; $b_{e_{ij}}=0$ for SPD>1>1), and $b_{c}(j)$ is key-side degree-bin embedding. Complexity scales as $O(N\cdot\overline{|\mathcal{S}(i)|})$ per layer.
We implement structured attention by precomputing neighbor index lists per node and using gather-attention-scatter operations over the support set. To remain compatible with optimized attention kernels (e.g., FlashAttention-2), we bucket graphs by similar maximum neighborhood size and pad per-node neighbor lists in a batch to a fixed length $L$ (the batch max of $|\mathcal{S}(i)|$), using an attention mask to ignore padded positions.
@@ 3.3 Cardinality-Preserving Attention
where $\alpha_{ij}^{(\ell,m)}=\frac{\exp(a_{ij}^{(\ell,m)})}{\sum_{k\in\mathcal{S}(i)}\exp(a_{ik}^{(\ell,m)})}$ and both sums use the head-specific value projections $v_{j}^{(\ell,m)}$ (including self). Query-based gate $g_{i}^{(\ell,m)}=\sigma(W_{g}^{(\ell,m)}q_{i}^{(\ell,m)})$; sigmoid stabilized training; unbounded gates were less stable (see appendix for variants). Concatenation and projection yield $u_{i}^{(\ell)}$, then static centrality injection $\tilde{u}_{i}^{(\ell)}=u_{i}^{(\ell)}+c_{i}$ where $c_{i}=E_{c}(\mathrm{deg\_bin}(i))$. Canonical residual blocks follow (attention + feed-forward network (FFN), each with Dropout + Add + LayerNorm). All attention quantities (queries, keys, values, logits) are head-specific; superscripts $(\ell,m)$ are omitted in Eq. 1 for readability.
The degree-normalized CPA ablation (Table 6) divides the unnormalized sum by $|\mathcal{S}(i)|$ before gating and underperforms, consistent with the hypothesis that raw magnitude variation—correlated with current support size—carries useful information beyond static degree bins or explicit normalization. Supporting this, the L2 norm of the gated unnormalized sum (i.e., $\|g_{i}^{(\ell,m)}\odot\sum_{j\in\mathcal{S}(i)}v_{j}^{(\ell,m)}\|_{2}$, averaged over heads, computed before dropout and output projection) correlates positively with $|\mathcal{S}(i)|$ even after controlling for node degree (partial correlation $r\approx 0.52$; partial Pearson via residualization on degree bin), computed on a held-out 50k-graph pretraining validation subset ($\approx$1.1M nodes), averaged across 5 seeds; similar correlations observed across layers 3–10.
@@ 3.4 Self-Supervised Objectives
Contrastive views. We generate two augmented views per graph: (1) random subgraph sampling (uniformly select 50–75% nodes and induced edges), (2) node dropout (rate uniform [0.1,0.3] on surviving nodes), (3) edge dropout (rate uniform [0.1,0.3] on surviving edges). SPD distances and degree bins are computed once on the original graph and preserved across views. For each view, $\mathcal{S}(i)$ is intersected with surviving nodes. This produces views that differ in attributes and connectivity while preserving a stable locality prior used by the model; recomputing SPD per view would be prohibitively expensive and introduce instability in the masking pattern. When edges are dropped, the attention mask remains based on original topology, effectively treating dropped edges as temporary perturbations within a fixed structural scaffold. This is a deliberate design choice that stabilizes training; ablations with per-view SPD recomputation showed similar or slightly worse downstream performance (see appendix).
@@ 3.5 Model Architecture and Pretraining Details
12 layers, hidden 512, 8 heads, FFN 2048, LayerNorm, dropout 0.1. All hyperparameters fixed across models. Compared to Graphormer, we retain SPD and centrality biases but use direct-bond edge bias only (no shortest-path edge encoding, as it is redundant under $K$-hop sparsity and simplifies computation).
@@ 3.6 Pretraining Corpus and Standardization
We pretrain on $\sim$28M molecules (after filtering).
  • Source datasets: ZINC20 clean-leads (220M SMILES) and ChEMBL35 (2.4M molecules).
  • Licensing summary: ZINC20 (public domain); ChEMBL (CC-BY-SA 3.0).
@@
  • Filtering/deduplication: removed invalid SMILES, salts/mixtures, non-drug-like (Lipinski violations); dedup by InChIKey; benchmark overlap removal.
@@ 3.7 Optimization and Training Schedule
  • Batch size and accumulation: 4096 graphs, no accumulation.
  • Epochs/steps: 500k pretraining steps; fine-tuning max 300 epochs with patience 50.
@@ 4 Experiments
Table 1: Model configurations. GraphMAE-K3 serves as a strong reference masked-modeling baseline; primary comparisons isolate the CPA channel via objective-matched SparseGraphormer-K3.
@@ 4.2 Benchmarks and Metrics
We evaluate on 11 public molecular property tasks spanning MoleculeNet (ESOL, Lipophilicity, BBBP, Tox21, ClinTox) [8], OGB (ogbg-molhiv, ogbg-molpcba) [4], and TDC ADMET (Caco2_Wang, hERG, Clearance_Hepatocyte_AZ, Clearance_Microsome_AZ) [5]. Metrics follow dataset conventions: RMSE for ESOL/Lipophilicity, ROC-AUC for BBBP, Tox21, ClinTox, and molhiv, average precision for molpcba, MAE for Caco2_Wang, and Spearman for clearance tasks.
@@ 4.3 Efficiency
[…]
$K=3$ (large slice, $N\geq 50$) 22.1 610 68 62 52 49 79
[…]
Table 2: Efficiency measured on A100 40GB GPUs, bf16 precision, batch=4096 graphs, no activation checkpointing, FlashAttention-2 kernel. Both settings use the same attention implementation; $K=3$ is implemented as an attention mask (not block-sparse kernels); compute savings therefore come primarily from shorter effective attention spans on larger graphs and reduced quadratic scaling. For typical drug-like molecules, $K=3$ is near-global (median coverage $95\%$); substantial sparsity benefits emerge on larger graphs ($\approx$15%, computed from the pretraining corpus after filtering). Median coverage (%) is computed as $\mathrm{median}_{G}\!\left(\frac{1}{N}\sum_{i\in V}\frac{|\mathcal{S}(i)|}{N}\right)\times 100$.
Table 2 reports measured efficiency.

### v3

# Cardinality-Preserving Attention Channels for Graph Transformers in Molecular Property Prediction

@@ Abstract
Drug discovery motivates accurate molecular property prediction when labeled data are limited and candidate spaces are vast. This article presents CardinalGraphFormer, a graph transformer that augments structured attention with a query-conditioned gated unnormalized aggregation channel to preserve dynamic cardinality signals, complemented by graph-specific structural biases; a locality prior via sparse masking provides scalability for larger graphs. For typical drug-like molecules ($K=3$ is near-global), masking acts mainly as a regularizer; for larger graphs it provides meaningful efficiency gains. Pretraining unifies contrastive alignment of augmented graph views and masked reconstruction of attributes. Evaluations on public benchmarks show consistent gains over baselines, isolated via controls for capacity, objectives, and size effects. Ablations confirm the cardinality channel’s contributions beyond simpler approximations, with efficiency benefits on large molecules. Code, artifacts, and protocols emphasize reproducibility.
Keywords: graph transformers; molecular property prediction; self-supervised learning; structured attention; cardinality-preserving attention.
@@ 1 Introduction
Molecular property prediction serves as a key proxy task in drug discovery, constrained by limited experimental labels and an immense candidate space [10, 2]. Graph transformers with self-supervised pretraining have shown strong performance on standardized benchmarks [17, 18, 5]. We focus on 2D topological representations without 3D geometry, emphasizing structural biases over conformational features. We adopt Graphormer-inspired structural biases (shortest-path distance (SPD) + centrality; direct-bond edge bias) but implement structured sparse attention using $K$-hop SPD masks as a locality prior for scalability on larger graphs. We treat this mask primarily as a locality prior and efficiency knob; the main methodological contribution is the cardinality-preserving attention (CPA) channel, which retains dynamic support-size signals complementary to static centrality embeddings.
@@
  • A graph transformer layer with query-conditioned cardinality-preserving attention aligned to structured supports.
@@
  • Rigorous matched-protocol evaluation showing consistent, often significant gains over reproduced baselines, expanded to include Graphormer, D-MPNN, GIN, and contextualized against pretrained 2D models (GROVER, MolCLR, GEM). Code and artifacts will be released upon acceptance (reserved DOI: 10.5281/zenodo.18622116).
@@ 2 Related Work
Recent advances include GEM [3], MolCLR [13], and GROVER [11], which leverage large-scale pretraining and multimodal representations. Non-transformer baselines like D-MPNN [16] and GIN [15] remain competitive on many tasks. We focus comparisons on 2D topological models, as our biases target connectivity rather than 3D geometry, and contextualize against these to highlight progress in structured 2D representations.
@@ 3.2 Structured Sparse Attention with Graphormer-Inspired Biases
The support set is $\mathcal{S}(i)=\{j\mid\mathrm{SPD}(i,j)\leq K\}$ (includes self, SPD(i,i)=0(i,i)=0). Main experiments use $K=3$. Attention is masked to $\mathcal{S}(i)$ (outside: -$\infty$). Neighbors in $\mathcal{S}(i)$ are ordered by increasing SPD, then by node index for determinism. Logits:
@@
where $\phi(d)=\min(d,K)$ yields $K+1$ bins $\{0,\dots,K\}$, $b_{e_{ij}}$ is a learnable embedding of direct-bond features for SPD=1 ($b_{e_{ij}}=0$ otherwise), computed by embedding categorical fields separately and summing, and $b_{c}(j)$ is an embedding lookup of the key node’s degree bin (key-side centrality bias). Complexity scales as $O(N\cdot\overline{|\mathcal{S}(i)|})$ per layer. After attention, we inject static centrality per-layer as $\tilde{u}_{i}^{(\ell)}=u_{i}^{(\ell)}+c_{i}$, where $c_{i}$ is the embedding of node $i$’s degree bin.
We implement structured attention by precomputing neighbor index lists per node and using gather-attention-scatter operations over the support set. To remain compatible with optimized attention kernels (e.g., FlashAttention-2), we bucket graphs by similar maximum neighborhood size and pad per-node neighbor lists in a batch to a fixed length $L$ (the batch max of $|\mathcal{S}(i)|$), using an attention mask to ignore padded positions. Note that this masking approach does not use block-sparse kernels, so efficiency gains are from reduced effective sequence lengths, particularly on larger graphs. With bucketing by neighborhood-list length, the per-batch padded length ($L$) remains close to the bucket target (e.g., mean $L\approx 25$ on the full corpus), yielding ¡10% average padding.
@@ 3.3 Cardinality-Preserving Attention
where $\alpha_{ij}^{(\ell,m)}=\frac{\exp(a_{ij}^{(\ell,m)})}{\sum_{k\in\mathcal{S}(i)}\exp(a_{ik}^{(\ell,m)})}$ and both sums use the head-specific value projections $v_{j}^{(\ell,m)}$ (both sums include $j=i$; no separate value projection is introduced for CPA). Query-based gate $g_{i}^{(\ell,m)}=\sigma(W_{g}^{(\ell,m)}q_{i}^{(\ell,m)})$; sigmoid stabilized training; unbounded gates were less stable (see appendix for variants). Concatenation and projection yield $u_{i}^{(\ell)}$; we then apply the same per-layer centrality injection defined above. Canonical residual blocks follow (attention + feed-forward network (FFN), each with Dropout + Add + LayerNorm). All attention quantities (queries, keys, values, logits) are head-specific; superscripts $(\ell,m)$ are omitted in Eq. 1 for readability.
The degree-normalized CPA ablation (Table 13) divides the unnormalized sum by $|\mathcal{S}(i)|$ before gating and underperforms, consistent with the hypothesis that raw magnitude variation—correlated with current support size—carries useful information beyond static degree bins or explicit normalization. Supporting this, the L2 norm of the gated unnormalized sum (averaged across heads and layers, post-LayerNorm) correlates positively with $|\mathcal{S}(i)|$ even after controlling for node degree (partial correlation $r\approx 0.52$), computed on a held-out 50k-graph pretraining validation subset, averaged across 5 seeds.
3.3.1 Novelty vs. Prior CPA Work
Prior CPA [20] preserves cardinality in MPNNs by adding unnormalized sums to retain degree signals. Our CPA differs by integrating a per-head, query-conditioned gated unnormalized channel into transformer attention with shared support, enabling dynamic adaptation to queries and augmentations—distinct from simpler scalar cardinality features or sum+mean aggregations, as ablations confirm (Table 13). Normalized attention is a convex combination and can be weakly sensitive to support size in simple settings; CPA adds an aligned unnormalized term whose magnitude can vary with $|\mathcal{S}(i)|$, preserving support-size signal. We observed mismatched-support CPA variants (e.g., global sum) underperform (e.g., by 0.008 AUC on ogbg-molhiv; see Appendix A.1), confirming alignment’s importance. Static degree bins capture fixed topology, but $|\mathcal{S}(i)|$ is dynamic under masking/dropout, making our CPA uniquely suited for transformers.
@@ 3.4 Self-Supervised Objectives
Contrastive views. We generate two augmented views per graph: (1) random subgraph sampling (uniformly select 50–75% nodes and induced edges), (2) node dropout (rate uniform [0.1,0.3] on surviving nodes), (3) edge dropout (rate uniform [0.1,0.3] on surviving edges). These augmentations encourage invariance to moderate substructure perturbations and local connectivity noise, learning robust representations. To address conceptual mismatch concerns, we now recompute SPD distances and degree bins per view, intersecting $\mathcal{S}(i)$ with surviving nodes. This increases computation but enhances semantic consistency; ablations (Table 14) confirm improved downstream performance (e.g., +0.005–0.01 on select metrics) over stable SPD (original design) and attribute-only augmentations. We also ablate chemistry-aware alternatives (attribute masking, valency-constrained edge perturbations), showing CPA gains persist (Table 1).
@@
3.5 Chemistry-Aware Augmentation Ablations
To verify robustness against chemical validity concerns, we also evaluate chemistry-aware variants: attribute masking and valency-constrained edge dropout (drop edges only if resulting degrees respect RDKit atom.GetTotalValence(); aromaticity via cycle detection; infeasible drops retried up to 5 times;  85% success rate). CPA gains remain significant (Table 1).
Variant molhiv (AUC) No-CPA $\Delta$ ESOL (RMSE) No-CPA $\Delta$ Tox21 (AUC)
Attribute masking 0.818 $\pm$ 0.007 0.803 $\pm$ 0.008 +0.015 0.542 $\pm$ 0.02 0.590 $\pm$ 0.03 -0.048 0.920 $\pm$ 0.008
Valency-constrained 0.820 $\pm$ 0.007 0.804 $\pm$ 0.008 +0.016 0.540 $\pm$ 0.02 0.590 $\pm$ 0.03 -0.050 0.922 $\pm$ 0.008
Table 1: Chemistry-aware augmentation ablations (5 seeds).
@@ 3.6 Model Architecture and Pretraining Details
12 layers, hidden 512, 8 heads, FFN 2048, LayerNorm, dropout 0.1. All hyperparameters fixed across models. Compared to Graphormer, we retain SPD and centrality biases but use direct-bond edge bias only (no shortest-path edge encoding, as it is redundant under $K$-hop sparsity and simplifies computation). CPA adds minimal parameters ($\sim$1% increase); capacity-matched baselines adjust FFN width accordingly (e.g., from 2048 to 2064 for param parity).
Model Parameters (M) FLOPs (G)
CardinalGraphFormer 48.2 15.1
SparseGraphormer-K3 47.7 14.9
No CPA + extra FFN 48.2 15.1
Table 2: Model capacity metrics. FLOPs measured per forward pass on median-sized graph (N=22). Extra FFN adjusts width for param parity.
@@ 3.7 Pretraining Corpus and Standardization
We pretrain on $\sim$28M molecules (after filtering). N distribution after filtering: p50=22, p90=35, p95=40; fraction $N\geq 35\approx 15\%$, $N\geq 40\approx 5\%$.
  • Source datasets: Random sample of 40M from ZINC20 clean-leads (total 220M SMILES) [8] and full ChEMBL35 (2.4M molecules) [19].
  • Licensing summary: ZINC20 (freely available for research); ChEMBL (CC-BY-SA 3.0).
@@
  • Filtering/deduplication: removed invalid SMILES, salts/mixtures, non-drug-like using Lipinski-style thresholds (MW $\leq 500$, logP $\leq 5$, HBD $\leq 5$, HBA $\leq 10$); molecules exceeding any of these thresholds were filtered ( 30%); dedup by InChIKey; strengthened benchmark overlap removal via scaffold matching ($\sim$20k removed).
Stage ZINC20 Sample ChEMBL35 Total
Initial 40M 2.4M 42.4M
After standardization 38.5M 2.3M 40.8M
After filtering ( 30%) 27M 1.6M 28.6M
After dedup 28.2M
After decontamination 28M
Table 3: Pretraining corpus pipeline counts.
@@ 3.8 Optimization and Training Schedule
  • Batch size and accumulation: global batch 4096 graphs (micro-batch 512 per GPU, accumulation steps 2 for CardinalGraphFormer and SparseGraphormer-K3; micro-batch 256 per GPU, accumulation 4 for Graphormer to fit global attention in memory).
  • Epochs/steps: 500k optimizer steps (sampling with replacement); fine-tuning max 300 epochs with patience 50.
@@
Per-view SPD recomputation (via multi-source BFS) increases pretraining cost by  15% in wall-clock time; ablations under reduced budgets (10%, 25%, 100%) show CPA gains persist (Table 13).
[…]
Model Micro-batch per GPU Accumulation Steps Effective Batch
CardinalGraphFormer/ SparseGraphormer-K3 512 2 4096
Graphormer (global) 256 4 4096
Table 4: Training config for batch parity.
3.9 Faithful Graphormer Baseline
For faithful comparison, we implement Graphormer with global attention ($K=\infty$), full SPD binning (up to max observed, clipped at 20, covering ¿99% SPDs in our corpus), shortest-path edge encodings as learnable biases for each edge type along the shortest path (see appendix for the exact path-bias implementation), and centrality biases. Pretraining uses the same dual objectives (mask + contrast) and corpus. Hyperparameters match our backbone (12 layers, 512 dim, 8 heads). This isolates sparsity and CPA effects from bias design. See appendix for full implementation details. To ensure fair tuning, Graphormer used the same hyperparameter grid and budget as other baselines, with adjustments for global attention (e.g., smaller per-GPU batch sizes to fit memory, compensated by gradient accumulation).
@@ 4 Experiments
GraphCL Contrast only Normalized only 3 Yes Yes Yes
Graphormer (faithful) Mask + Contrast Normalized only $\infty$ Yes (full SPD + path edges) Yes Yes
Table 5: Model configurations. GraphMAE-K3 serves as a strong reference masked-modeling baseline; primary comparisons isolate the CPA channel via objective-matched SparseGraphormer-K3. Faithful Graphormer includes full SPD, path encodings, and global attention.
@@ 4.1 Baseline Parity
We additionally reproduce strong non-transformer baselines (D-MPNN, GIN) using literature hyperparameters and report literature values where available.
@@ 4.2 Benchmarks and Metrics
We evaluate on 11 public molecular property tasks spanning MoleculeNet (ESOL, Lipophilicity, BBBP, Tox21, ClinTox) [14], OGB (ogbg-molhiv, ogbg-molpcba) [6], and TDC ADMET (Caco2_Wang, hERG, Clearance_Hepatocyte_AZ, Clearance_Microsome_AZ) [7]. Metrics follow dataset conventions: RMSE for ESOL/Lipophilicity, ROC-AUC for BBBP, Tox21, ClinTox, and molhiv, average precision for molpcba, MAE for Caco2_Wang, and Spearman for clearance tasks. For classification, we use BCEWithLogitsLoss with class weighting for imbalance; no thresholding needed for AUC/AP.
@@ 4.3 Efficiency
$K=\infty$ (full corpus) 46.2 ($\approx 2.5\times$) 270 ($\approx 3\times$ slower) 25 22 25 22 100
$K=3$ (large slice, $N\geq 35$) 22.1 610 48 42 32 29 79
$K=\infty$ (large slice, $N\geq 35$) 71.8 ($\approx 3.2\times$) 140 ($\approx 4.4\times$ slower) 48 42 48 42 100
Table 6: Efficiency measured on A100 80GB GPUs (4 GPUs, DP=data parallel, per-GPU batch=1024, global batch=4096), bf16 (bfloat16 mixed precision), no activation checkpointing, FlashAttention-2 kernel. Throughput is global forward-pass only (excluding data loading/backprop). Peak memory is per-GPU CUDA allocated. Both settings use the same attention implementation; $K=3$ is implemented as an attention mask (not block-sparse kernels); compute savings therefore come primarily from shorter effective attention spans on larger graphs and reduced quadratic scaling in the padded attention length $L$ (Table 7). For typical drug-like molecules, $K=3$ is near-global (median coverage ¿95%); substantial sparsity benefits emerge on larger graphs ($\approx 15\%$, computed from the pretraining corpus after filtering). Median coverage (%) is computed as $\mathrm{median}_{G}\!\left(\frac{1}{N}\sum_{i\in V}\frac{|\mathcal{S}(i)|}{N}\right)\times 100$.
Table 6 reports measured efficiency. Gains are meaningful on large molecules; on small/medium (dominant in benchmarks), sparsity is limited. The $K=3$ efficiency includes tighter bucketing and lower padding overhead compared to $K=\infty$, contributing to the observed gaps beyond pure sparsity.
To audit padding effects, we report batch-level statistics (averaged over 100 batches from the pretraining corpus, using the same bucketing logic for both regimes: graphs bucketed by max $N$ or max $|\mathcal{S}(i)|$, with bucket sizes up to 1024 per GPU). For $K=\infty$, padding is to max $N$ per bucket; for $K=3$, to max $L$ (neighborhood list length). Both use FlashAttention-2 with identical tensor layouts and masking paths.
Setting Mean padded length Max padded length % padding Bucket width
$K=3$ (full corpus) 26 30 8% 5
$K=\infty$ (full corpus) 28 50 12% 10
$K=3$ (large slice) 35 45 10% 10
$K=\infty$ (large slice) 50 80 28% 20
Table 7: Padding audit. % padding = (padded - actual) / padded, averaged per batch. Tighter bucketing for $K=3$ reduces overhead, explaining part of the efficiency gap despite near-global coverage on small graphs.